# RBI-ThPhys-2025-26

### v1

@@ 4.1 One shape field
We begin by considering multipoles of $B_{\delta\delta E}$ in the top line of Eq. (49). As it happens, the lowest order multipole $\tilde{B}_{\delta\delta E}^{00}$ does not depend on $b_{2,3}^{\text{g}}$. Therefore, in order to constrain all bias parameters we need to include higher order multipoles as well. The degeneracy is broken extremely efficiently by the two quadrupoles $\tilde{B}_{\delta\delta E}^{20},\tilde{B}_{\delta\delta E}^{11}$, and Figure 6 illustrates this.
@@
The performance metrics discussed in Section 3 are shown in Figure 7, both for the Gaussian covariance $C_{\text{gauss}}$ and $\hat{C}_{\text{samp}}$ from Eq. (32). We observe good consistency in the FoB between the two choices. However, in the case of $\chi_{\text{red}}^{2}$ we see that the Gaussian covariance yields systematically lower values. This does not impact the reach of the model much according to the $\Delta(k_{\text{max}})$ criterion, as seen by examining the two blue lines in Figure 4. Moreover, we also checked that for $k_{\text{max}}\gtrsim 0.14\,h/\text{Mpc}$ some of the bias parameters change values significantly, which is a sign of overfitting since the values of the bias parameters should be uniquely determined. Altogether, we conclude that the three lowest order multipoles of $\tilde{B}_{\delta\delta E}$ are well fit by the tree-level EFT of IA up to at least $k_{\text{max}}\sim 0.11\,h/\text{Mpc}$ (and likely somewhat beyond) and in good agreement with the power spectrum.
Figure 9: Same as Figure 8 but now for multipoles with two shape fields.
@@
Lastly, we also investigated the monopole of $B_{\delta\delta B}$ and observe that it is in good agreement with what is expected from $\tilde{B}_{\delta\delta E}$; note that $B_{\delta\delta B}^{00}$ only depends on $b_{2,1}^{\text{g}}$ as mentioned in Paper I. In all our plots of the measured signal, we show error bars corresponding to the total volume $V_{\text{DESI}}=7.5\,(\text{Gpc}/h)^{3}$ of the DESI LRG sample [^14] Note that this is 3$\sim$ 3$\times$ 3 × larger than the volume considered in Paper I, where we assumed a smaller sky fraction due to possibly limited overlap with LSST.. More concretely, Figure 8 shows the measured bispectrum signal [^15] Note that in the equilateral (middle) panel, both the signal and the error bar are exactly zero (that is, to machine precision) for odd bispectra, due to symmetry in the equilateral configuration. To see this, note that if bin 1 and bin 2 are the same and $X=Y$ in Eq. (27), swapping $\mathbf{q}_{1}$ and $\mathbf{q}_{2}$ in the sum leaves the summand invariant, but the cross product acquires a minus sign, so that the estimator equals minus itself, and hence equals zero. Thus, the covariance is also zero in this case. for all four multipoles considered in this subsection, with the error bars rescaled to $V_{\text{DESI}}=7.5\,(\text{Gpc}/h)^{3}$. We also overplot the theory curves using the parameter choices
$b_{1}^{\text{g}}$ $=-0.070,$ (51)
$b_{2,1}^{\text{g}}$ $=0.015,$
$b_{2,2}^{\text{g}}$ $=0.12,$
$b_{2,3}^{\text{g}}$ $=-0.045.$
@@
Even for this larger volume, the model is broadly in agreement with the data. However, one can also see that the equilateral $\tilde{B}_{\delta\delta E}$ bispectrum does not match beyond $k\sim 0.10\,h/\text{Mpc}$ for $(\ell_{1},\ell_{2})=(0,0)$ and $(\ell_{1},\ell_{2})=(1,1)$ and ‘folded’ configurations of the form $(k,k/2,k/2)$ are also poorly fit (this is also the region where $\tilde{B}_{\delta\delta E}$ has the highest SNR, so it is perhaps not surprising that the model fails there sooner). Observe that the parity-odd multipole $B_{\delta\delta B}^{00}$ is clearly nonzero for the obtuse configurations, but vanishes for equilateral triangles because of symmetries (see Figure 2 in Paper I).
@@ 4.2 One or two shape fields
For bispectra with two tensor fields, to wit $B_{\delta EE},B_{\delta EB}$ and $B_{\delta BB}$, an additional parameter enters the tree-level model. This contribution was first derived in (31) and labelled as $\alpha_{\text{gg}}^{1\delta}$ in Paper I.
@@
Figure 11: Same as Figure 8 but now for multipoles with three shape fields. Observe again that the odd bispectra vanish in the equilateral configuration (Figure 6 of Paper I) because of symmetries (and the fact that two of the fields are the same).
@@
Finally, as a means of checking the compatibility with the parity-odd multipole $B_{\delta EB}^{02}$ again we also show in Figure 9 the bispectrum signal plotted for the set of parameters listed in Eq. (51) (that is, the same set that was used for the bispectra with one tensor field) along with the additional choice of $\alpha_{\text{gg}}^{1\delta}=4$. These bispectra are generally noisier than the ones with only one shape field, so the agreement is visually better. Note that this time, the parity-odd bispectrum $B_{\delta EB}$ does have a strong signal in the equilateral configuration (see also Figure 4 in Paper I); it no longer vanishes because of symmetries since all three fields are distinct.
@@ 4.3 One, two or three shape fields
$b_{\delta K}$ $=\frac{2}{3}b_{2,1}^{\text{g}}+\frac{1}{3}b_{2,3}^{\text{g}}=0.000\pm 0.007;$ (54)
$b_{K^{2}}$ $=b_{2,1}^{\text{g}}-b_{2,3}^{\text{g}}=0.060\pm 0.021;$
$b_{t}$ $=\frac{7}{8}(b_{2,3}^{\text{g}}-b_{2,2}^{\text{g}})=0.131\pm 0.021;$
@@ Acknowledgements.
We thank Fabian Schmidt for useful discussions throughout this project. This publication is part of the project “A rising tide: Galaxy intrinsic alignments as a new probe of cosmology and galaxy evolution” (with project number VI.Vidi.203.011) of the Talent programme Vidi which is (partly) financed by the Dutch Research Council (NWO). Z.V. acknowledges the support of the Kavli Foundation.
@@ Appendix A Theory Expressions for Tree-level Tensor Bispectrum
$B_{ijklrs}^{\alpha\beta\gamma,\text{det,tree}}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})$ $=2\,\mathcal{K}_{ij}^{\alpha,(1)}(\mathbf{k}_{1})\mathcal{K}_{kl}^{\beta,(1)}(\mathbf{k}_{2})\mathcal{K}_{rs}^{\gamma,(2)}(\mathbf{k}_{1},\mathbf{k}_{2})P_{L}(k_{1})P_{L}(k_{2})+\text{2 perm.}$. (57)
@@
[…]
@@
$B_{GGE}$$\boldsymbol{\times}$
$B_{GGB}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
@@
$B_{GBB}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
$B_{GEB}$$\boldsymbol{\times}$
@@
$B_{EEB}$$\boldsymbol{\times}$
$B_{EBB}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$

### v2

@@ 4.1 One shape field
We begin by considering multipoles of $B_{\delta\delta E}$ in the top line of Eq. (49). The lowest order multipole $\tilde{B}_{\delta\delta E}^{00}$ does not depend on $b_{2,3}^{\text{g}}$. Therefore, in order to constrain all bias parameters we need to include higher order multipoles as well. The degeneracy is broken extremely efficiently by the two quadrupoles $\tilde{B}_{\delta\delta E}^{20},\tilde{B}_{\delta\delta E}^{11}$, and Figure 6 illustrates this.
@@
The performance metrics discussed in Section 3 are shown in Figure 7, both for the Gaussian covariance $C_{\text{gauss}}$ and $\hat{C}_{\text{samp}}$ from Eq. (32). We observe good consistency in the FoB between the two choices. However, in the case of $\chi_{\text{red}}^{2}$ we see that the Gaussian covariance yields systematically lower values. This does not impact the reach of the model much according to the $\Delta(k_{\text{max}})$ criterion, as seen by examining the two blue lines in Figure 4. Moreover, we also checked that for $k_{\text{max}}\gtrsim 0.14\,h/\text{Mpc}$ some of the bias parameters change values significantly, which is a sign of overfitting since the values of the bias parameters should be uniquely determined. Altogether, we conclude that the three lowest order multipoles of $B_{\delta\delta E}$ are well fit by the tree-level EFT of IA up to at least $k_{\text{max}}\sim 0.11\,h/\text{Mpc}$ (and likely somewhat beyond) and in good agreement with the power spectrum.
Figure 9: Same as Figure 8 but now for multipoles with two shape fields. The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.
@@
Lastly, we also investigated the monopole of $B_{\delta\delta B}$ and observe that it is in good agreement with what is expected from $B_{\delta\delta E}$; note that $B_{\delta\delta B}^{00}$ only depends on $b_{2,1}^{\text{g}}$ as mentioned in Paper I. In all our plots of the measured signal, we show error bars corresponding to the total volume $V_{\text{DESI}}=7.5\,(\text{Gpc}/h)^{3}$ of the DESI LRG sample [^14] Note that this is $\sim 3\times$ larger than the volume considered in Paper I, where we assumed a smaller sky fraction due to possibly limited overlap with LSST.. More concretely, Figure 8 shows the measured bispectrum signal [^15] Note that in the equilateral (middle) panel, both the signal and the error bar are exactly zero (that is, to machine precision) for odd bispectra, due to symmetry in the equilateral configuration. To see this, note that if bin 1 and bin 2 are the same and $X=Y$ in Eq. (27), swapping $\mathbf{q}_{1}$ and $\mathbf{q}_{2}$ in the sum leaves the summand invariant, but the cross product acquires a minus sign, so that the estimator equals minus itself, and hence equals zero. Thus, the covariance is also zero in this case. for all four multipoles considered in this subsection, with the error bars rescaled to $V_{\text{DESI}}=7.5\,(\text{Gpc}/h)^{3}$. We also overplot the theory curves using the parameter choices
$b_{1}^{\text{g}}$ $0,\displaystyle=-07$0, (51)
$b_{2,1}^{\text{g}}$ $0,\displaystyle=02$0,
[…]
$b_{2,3}^{\text{g}}$ $0.\displaystyle=-04$0.
@@
These are the best-fit model parameters from the joint fit of all multipoles (cf. Section 4.3). Even for this larger volume, the model is broadly in agreement with the data. However, one can also see that the equilateral $\tilde{B}_{\delta\delta E}$ bispectrum does not match beyond $k\sim 0.10\,h/\text{Mpc}$ for $(\ell_{1},\ell_{2})=(0,0)$ and $(\ell_{1},\ell_{2})=(1,1)$ and ‘folded’ configurations of the form $(k,k/2,k/2)$ are also poorly fit (this is also the region where $\tilde{B}_{\delta\delta E}$ has the highest SNR, so it is perhaps not surprising that the model fails there sooner). Observe that the parity-odd multipole $B_{\delta\delta B}^{00}$ is clearly nonzero for the obtuse configurations, but vanishes for equilateral triangles because of symmetries (see Figure 2 in Paper I).
@@ 4.2 One or two shape fields
For bispectra with two tensor fields, i.e. $B_{\delta EE},B_{\delta EB}$ and $B_{\delta BB}$, an additional parameter enters the tree-level model. This contribution was first derived in (31) and labelled as $\alpha_{\text{gg}}^{1\delta}$ in Paper I.
@@
Figure 11: Same as Figure 8 but now for multipoles with three shape fields. Observe again that the odd bispectra vanish in the equilateral configuration (Figure 6 of Paper I) because of symmetries (and the fact that two of the fields are the same). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.
@@
Finally, as a means of checking the compatibility with the parity-odd multipole $B_{\delta EB}^{02}$ again we also show in Figure 9 the bispectrum signal plotted for the set of parameters listed in Eq. (51) (that is, the same set that was used for the bispectra with one tensor field) along with the additional choice of $\alpha_{\text{gg}}^{1\delta}=4.5$ (cf. Eq. (53)). These bispectra are generally noisier than the ones with only one shape field, so the agreement is visually better in most cases. However, the amplitude of $B_{\delta BB}^{02}$ prefers lower values of $\alpha_{\text{gg}}^{1\delta}$ than those preferred by $B_{\delta EE}^{02}$, as exemplified by the middle left plot. This is simply due to the latter dominating the overall SNR. Note that this time, the parity-odd bispectrum $B_{\delta EB}$ does have a strong signal in the equilateral configuration (see also Figure 4 in Paper I); it no longer vanishes because of symmetries since all three fields are distinct.
@@ 4.3 One, two or three shape fields
[…]
@@ Acknowledgements.
We thank Fabian Schmidt for useful discussions. This publication is part of the project “A rising tide: Galaxy intrinsic alignments as a new probe of cosmology and galaxy evolution” (with project number VI.Vidi.203.011) of the Talent programme Vidi which is (partly) financed by the Dutch Research Council (NWO). Z.V. acknowledges the support of the Kavli Foundation.
@@ Appendix A Theory Expressions for Tree-level Tensor Bispectrum
$B_{ijklrs}^{\alpha\beta\gamma,\text{det,tree}}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})$ $=2\,\mathcal{K}_{ij}^{\alpha,(1)}(\mathbf{k}_{1})\mathcal{K}_{kl}^{\beta,(1)}(\mathbf{k}_{2})\mathcal{K}_{rs}^{\gamma,(2)}(\mathbf{k}_{1},\mathbf{k}_{2})P_{L}(k_{1})P_{L}(k_{2})+\text{2 perm.}$ (57)
@@
$g_{ij}^{\text{stoch}}(\mathbf{x})\qquad\supset\qquad\epsilon_{ij}(\mathbf{x})\text{ (first order)};\qquad\epsilon_{ij}^{\delta}(\mathbf{x})\delta(\mathbf{x}),\quad\epsilon^{\Pi}(\mathbf{x})\text{TF}(\Pi^{[1]})_{ij}(\mathbf{x})\text{ (second order) }$ (64)
@@
[…]
@@
[…]
@@
[…]